A manufacturer sells his product at (49-2x) per unit and it costs him 2 shillings per unit to manufacture. if his fixed costs are 30 shillings
a)determine the break even point in terms of cost and quantity
b) determine the level output that will maximize profits
1
Expert's answer
2018-03-30T09:27:08-0400
A manufacturer sells his product at P = (49-2x) per unit and it costs him AVC = 2 shillings per unit to manufacture. If his fixed costs are FC = 30 shillings, then:
a) the break even point in terms of cost and quantity is FC/(P - AVC) = 30/(49 - 2x - 2) = 30/(47 - 2x).
b) the level output that will maximize profits is at MR = MC = P.
MC = AVC = 2, so:
49 - 2x = 2,
x = 23.5 units.